solve the equation for x belongs to R x/x-1 is equal to x + 1/x-2


1
Expert's answer
2021-12-16T17:02:07-0500

x/x-1 is equal to x + 1/x-2

"\frac{x}{x-1}=x+\frac{1}{x-2}\\
\frac{x}{x-1}=\frac{x^2-2x+1}{x-2}\\
\frac{x^2-2x+1}{x-2}-\frac{x}{x-1}=0\\
\frac{x^3-2x^2+x-x^2+2x-1-x^2+2x}{(x-2)(1-x)}=0\\
\frac{x^3-3x^2+5x-1}{(x-2)(1-x)}=0\\
x^3-3x^2+5x-1=0"

Solving the above equation by trial method, we get

"x=0.2290, 1.385,1.385"